OpenStudy (anonymous):

Express each of the following in terms of sinθ and/or cosθ. a) cosθ 1/secθ b) tanθcosθ c) 1+cot^2θ/cot^2θ

OpenStudy (anonymous):

\(\tan(x)=\frac{\sin(x)}{\cos(x)}\) so second one is \[\frac{\sin(x)}{\cos(x)}\times \cos(x)=\frac{\sin(x)}{\cancel{\cos(x)}}\times \cancel{\cos(x)}=\sin(x)\]

OpenStudy (anonymous):

third one \(1+cot^2(x)=\csc^2(x)\) so you get \[\frac{\csc^2(x)}{\cot^2(x)}\] \[=\frac{\frac{1}{\sin^2(x)}}{\frac{\cos^2(x)}{\sin^2(x)}}\] \[=\frac{1}{\sin^2(x)}\times \frac{\sin^2(x)}{\cos^2(x)}\] \[=\frac{1}{\cos^2(x)}\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Didn't get the first one